# On estimation of weighted cumulative residual Tsallis entropy

### v1

@@ 5 Applications
In this section, we develop a goodness-of-fit test for uniform distribution using an estimator of WCRTE measure. Uniform distribution is used for random number generation. Uniformity test is an important problem in statistics and this problem has been addressed by many researchers in the literature. For example, Stephens (1974) studied uniformity test using cdf-based statistic and Dudewicz and Van Der Meulen (1981) proposed an entropy-based test of uniformity. For recent development in this area, redears may refer to Noughabi (2022), Chakraborty et al. (2023) and the references therin.
@@
$n$ $\alpha$ $\mathcal{C}_{0.025,n}$ $\mathcal{C}_{0.975,n}$ $n$ $\alpha$ $\mathcal{C}_{0.05,n}$ $\mathcal{C}_{0.95,n}$
@@
5.1 Critical points and Power of the tests
@@ 6 Conclusion
In this paper, we study some new properties of WCRTE measure and develop four non-parametric estimators of WCRTE. We investigate asymptotic properties of the proposed estimators as well. Also, we propose estimators for WCRE measure and study their asymptotic properties. We compare the performances of these estimators using simulation and develop two new uniformity tests. Power of these tests are compared with some popular existing tests and it is observed that, proposed tests perform very well compared to the other tests in majority of the cases.
We use one estimator of WCRTE and WCRE to construct the uniformity test. Other estimators may also be used for this purpose. It will be interesting to see for which value of the window size of these estimators give highest power. More work is needed in this direction.

### v2

@@ 6 Applications
In this section, we develop a goodness-of-fit test for uniform distribution using an estimator of WCRTE measure for complete and progressively type-II censored data. Uniform distribution is used for random number generation. Uniformity test is an important problem in statistics and this problem has been addressed by many researchers in the literature. For example, Stephens (1974) studied uniformity test using cdf-based statistic and Dudewicz and Van Der Meulen (1981) proposed an entropy-based test of uniformity. For recent development in this area, redears may refer to Noughabi (2022), Chakraborty et al. (2023) and the references therin.
6.1 Uniformity Test for Complete Sample
@@
$n$ $\alpha$ $\mathcal{C}_{0.025,n}$ $\mathcal{C}_{0.975,n}$ $n$ $\alpha$ $\mathcal{C}_{0.025,n}$ $\mathcal{C}_{0.975,n}$
@@
6.1.1 Critical Points and Power of the Tests
@@
6.2 Uniformity under Progressively Type-II Censored Data
Now we extend the uniformity test for complete sample to progressively type-II censored sample. Let $X_{1:r:n},\cdots,X_{r:r:n}$ be a progressively type-II censored sample with censoring scheme $(R_{1},R_{2},\cdots,R_{r})$ from a continuous rv having density $f$ concentrated on [0,1]. Based on this sample, we want to test whether $f$ is uniformly distributed or not. We consider the estimator $\xi_{\alpha}^{w}P$ with $\alpha$ = 2. From Proposition 5.1, we found that $\xi_{\alpha}^{w}P$ is a consistent estimator when $r\to n$ i.e. $\xi_{\alpha}^{w}P$ reduces to $\hat{\xi}_{\alpha}^{w}(X)$ as $r\to n$. From Lemma 6.2 we have that $\hat{\xi}_{\alpha}^{w}(X)$ lies between (0,0.125) for $\alpha$ = 2. So, we can use $\xi_{\alpha}^{w}P$ as a test statistic for testing uniformity when the data are progressively type-II censored. Also note that, in Table 6 for U(0,1) all the average values of $\xi_{\alpha}^{w}P$ lie within the interval (0,0.125).
Table 9: Critical points of $\xi_{\alpha}^{w}P$.
$n$ $r$ Schemes $D_{0.025:r:n}$ $D_{0.975:r:n}$
20 10 (10,0910,0*9) 0.04620 0.10704
(09,100*9,10) 0.00712 0.04500
(5,08,55,0*8,5) 0.01471 0.07560
(1101*10) 0.01991 0.10209
15 (5,0145,0*14) 0.05614 0.10440
(014,50*14,5) 0.02711 0.08558
(07,5,070*7,5,0*7) 0.05259 0.10666
(3,013,23,0*13,2) 0.04077 0.09848
30 10 (20,0920,0*9) 0.04466 0.10752
(09,200*9,20) 0.00210 0.10866
(10,08,1010,0*8,10) 0.00632 0.04544
(2102*10) 0.00975 0.08349
20 (10,01910,0*19) 0.06153 0.10184
(019,100*19,10) 0.02377 0.07049
(5,018,55,0*18,5) 0.03758 0.08940
50 30 (20,02920,0*29) 0.06725 0.09926
(029,200*29,20) 0.02200 0.05540
(10,028,1010,0*28,10) 0.03750 0.08060
We reject the hypothesis if $\xi_{\alpha}^{w}P$ is large or small. The critical region for significance level $\gamma$ is
$\xi_{\alpha}^{w}P\leq D_{1-\frac{\gamma}{2}:r:n}\;\;\;\text{or}\;\;\;\xi_{\alpha}^{w}P\geq D_{\frac{\gamma}{2},r:n},$
where $D_{\gamma}$ is the $\gamma$-th quantile point of the empirical distribution of $\xi_{\alpha}^{w}P$. We simulate 5000 PC-II censored samples from U(0,1) distribution and calculate the critical points at 5% level of significance. The critical points are provided in Table 9. We compute power of the test for various progressive censoring schemes and for $\alpha$ = 2. We consider the same alternatives given by Stephens (1974). We generate 5000 PC-II censored samples from these alternatives and compute the power of the test and present them in Table 10. From the table we observe that the proposed test attains the nominal level of significance for all the censoring schemes. As the sample size increases the power also increases. In general, the performance of the proposed test is good for alternatives $A_{j}$’s and $C_{j}$’s. We have observed this phenomenon for complete data as well. Proposed test also performs well for alternative $B_{2}$. For alternatives $B_{1.5}$ and $B_{2}$, the test works well when sample size is high.
Table 10: Power of the test for various schemes for $\alpha$ = 2.
Alternatives
$n$ $r$ Schemes U(0,1) $A_{1.5}$ $A_{2}$ $B_{1.5}$ $B_{2}$ $B_{2}$ $C_{1.5}$ $C_{2}$
20 10 (10,0910,0*9) 0.054 0.154 0.373 0.097 0.214 0.556 0.109 0.201
(09,100*9,10) 0.053 0.185 0.488 0.013 0.006 0.010 0.263 0.496
(5,08,55,0*8,5) 0.049 0.168 0.433 0.023 0.030 0.096 0.175 0.370
(1101*10) 0.050 0.135 0.299 0.034 0.057 0.154 0.106 0.184
15 (5,0145,0*14) 0.048 0.207 0.553 0.137 0.352 0.818 0.130 0.304
(014,50*14,5) 0.051 0.223 0.578 0.057 0.123 0.494 0.171 0.374
(07,5,070*7,5,0*7) 0.047 0.183 0.465 0.107 0.296 0.709 0.114 0.241
(3,013,23,0*13,2) 0.050 0.214 0.570 0.097 0.257 0.710 0.152 0.314
30 10 (20,0920,0*9) 0.049 0.148 0.348 0.088 0.200 0.496 0.097 0.198
(09,200*9,20) 0.051 0.162 0.434 0.015 0.002 0.001 0.146 0.319
(10,08,1010,0*8,10) 0.048 0.167 0.437 0.008 0.004 0.004 0.231 0.464
(2102*10) 0.0.047 0.142 0.317 0.013 0.014 0.025 0.152 0.270
20 (10,01910,0*19) 0.052 0.290 0.714 0.180 0.519 0.947 0.183 0.424
(019,100*19,10) 0.050 0.322 0.762 0.046 0.124 0.529 0.255 0.544
(5,018,55,0*18,5) 0.053 0.324 0.750 0.110 0.335 0.832 0.185 0.418
50 30 (20,02920,0*29) 0.053 0.461 0.916 0.318 0.787 0.999 0.264 0.585
(029,200*29,20) 0.054 0.492 0.939 0.044 0.144 0.650 0.393 0.716
(10,028,1010,0*28,10) 0.049 0.470 0.920 0.150 0.481 0.956 0.256 0.596
6.3 Data Analysis
The problem of testing uniformity can be implemented in any general foodness-of-fit test. For example, consider the problem of testing
$H_{0}:X\sim f(X)\;\;\;\;vs.\;\;\;\;H_{1}:X\not\sim f(X).$
Using the transformation $Y=F(X)$, the above problem becomes
$H_{0}:Y\sim U(0,1)\;\;\;\;vs.\;\;\;\;H_{1}:Y\not\sim U(0,1).$
Now consider the failure times for insulating fluid dataset given in Subsection 4.1. This dataset follows exponential distribution with MLE, $\hat{\lambda}$ = 0.0696. Now, if we want to test exponentiality based on this dataset, then using the transformation $Y=1-\exp(-\hat{\lambda}X)$, the problem reduces to testing uniformity. The transformed data are as follows:
0.01314, 0.05284, 0.06463, 0.08714, 0.17592, 0.19743, 0.25087, 0.27750,
0.28650, 0.36390, 0.40044, 0.42736, 0.43763, 0.56802, 0.89028, 0.89600,
0.90559, 0.92231 0.99374.
Based on these 19 observations, the estimated WCRE value is 0.1534. Note that the true value of WCRE for U(0,1) distribution is 0.1389. The 5% critical points for WCRE-based uniformity test with $n$ = 19 are (0.09423, 0.15482). Since, the estimated test statistic lies within the critical points, we cannot reject the null hypothesis that $Y$ is uniformly distributed over [0,1]. This implies that we cannot reject the null hypothesis that $X$ is exponentially distributed. Similarly, using WCRTE we can perform uniformity test. For $\alpha$ = 2, the value of the estimated test statistic is 0.0880. The true value of WCRTE for U(0,1) is 0.0834. The 5% critical points are (0.05453, 0.09871). Since the estimated value of WCRTE lies between the critical points, we cannot reject the null hypothesis of uniformity of $Y$ over [0,1] and consequently exponentiality of $X$.
We now apply the proposed uniformity test for PC-II censored samples. We generate PC-II censored data from the above dataset with $r$ = 10. The generated observations and the associated censoring schemes are provided below.
  • Scheme 1: ($R_{1}=9,R_{2}=\cdots=R_{10}=0$).
    {0.0134, 0.05284, 0.06463, 0.19743, 0.25087, 0.28650, 0.40044, 0.42736, 0.56802, 0.89028}
  • Scheme 2: ($R_{1}=\cdots=R_{9}=0,R_{10}=9$).
    {0.01314, 0.05284, 0.06463, 0.08714, 0.17592, 0.19743, 0.25087, 0.27750, 0.28650, 0.36390}
  • Scheme 3: ($R_{1}=5,R_{2}=\cdots=R_{9}=0,R_{10}=4$).
    {0.01314, 0.05284, 0.06463, 0.08714, 0.17592, 0.25087, 0.27750, 0.40044, 0.42736, 0.43763}
  • Scheme 4: ($R_{1}=\cdots=R_{9}=1,R_{10}=0$).
    {0.01314, 0.05284, 0.06463, 0.08714, 0.17592, 0.19743, 0.25087, 0.36390, 0.40044, 0.89600}
We want to test the hypothesis that the PC-II samples come from U(0,1) distribution. We compute the test statistic and 5% critical points for every schemes and the obtained results are provided in Table 11. From the table we observe that the proposed test accepts the null hypothesis for all the schemes which is accurate.
Table 11: Uniformity test resluts for PC-II censoring.
Schemes Test Statistic Critical Points Decision
1 0.07115 (0.04477, 0.10707) Accept
2 0.01392 (0.00810, 0.05040) Accept
3 0.02196 (0.01771, 0.08343) Accept
4 0.09124 (0.02833, 0.11140) Accept
@@ 7 Conclusion
In this paper, we study some new properties of WCRTE measure and develop four non-parametric estimators of WCRTE. We investigate asymptotic properties of the proposed estimators as well. We propose estimators for WCRE measure and study their asymptotic properties. Also, we suggested an estimator for WCRTE under progressively type-II censored samples. We compare the performances of these estimators using simulation and develop two new uniformity tests for complete samples. Power of these tests are compared with some popular existing tests and it is observed that proposed tests perform better compared to the other tests in majority of the cases. Also, we proposed an uniformity test under progressive type-II censoring. We investigated its performance by evaluating the power and we have observed that the uniformity test for progressive type-II censoring works quite well.
We use one estimator of WCRTE and WCRE to construct the uniformity test. Other estimators may also be used for this purpose. It will be interesting to see, for which value of the window size of these estimators gives the highest power. We evaluate the power of the uniformity test under progressive type-II censoring when $\alpha$ = 2. Computation became very intensive when it came to progressive type-II censoring. Note that, for fixed values of $n$ and $r$, there exists $n-1\choose r-1$ number of different censoring schemes. Therefore the number of progressive censoring schemes becomes very large even for moderate $n$ and $r$. It will be interesting to study the power of the test for different choices of $\alpha$ and for other censoring schemes. More work is needed in this direction.